# Sorbonne Université & CNRS, 4 place Jussieu, 75252 Paris Cedex 05, France

### v1

@@ 3.1 Modelling signal processes at the LHC
Figure 2: Leading-order cross sections for single top production in association with a photon via $cg$ (or equivalently $\bar{c}g$) fusion (left), $\bar{u}g$ fusion (middle) and $ug$ fusion (right). The bands indicate theoretical uncertainties from scale and PDF variations summed in quadrature.
@@
as illustrated in figure 2 for a hadronic centre-of-mass energy $\sqrt{s}=13.6$ TeV. These predictions are obtained using Madgraph_aMC@NLO (v3.4.1) Alwall:2014hca , with the NNPDF_40_lo_as_01180 PDF set NNPDF:2021njg and renormalisation and factorisation scales set to the average transverse mass of the final-state particles, and an implementation of the relevant FCNC operators in FeynRules Christensen:2009jx ; Alloul:2013bka that has been exported in the UFO format Degrande:2011ua ; Darme:2023jdn . The resulting cross sections are found to span several orders of magnitude, from 1–10 fb for $\mathcal{O}(1)$ couplings down to 0.01–0.1 ab for couplings around $10^{-4}$ with the quadratic FCNC coupling dependence given in Eq. (11).
@@
This form for the cross section illustrates a clean separation between the QCD-driven $t\bar{t}$ production mechanism, one FCNC-induced top decay and one SM top decay. We compute the inclusive $t\bar{t}$ production cross section using Top++ v2.0 Czakon:2011xx ; Czakon:2013goa , including next-to-next-to-leading-order (NNLO) QCD corrections matched to next-to-next-to-leading-logarithmic (NNLL) soft-gluon resummation. We use this time the NNPDF_40_nnlo_as_01180 PDF set, and we fix both the renormalisation and factorisation scales to the top mass.
@@ 3.2 Technical setup
$\begin{split}pp\to t~{}(\to\ell^{+}\bar{\nu}_{\ell}b)~{}\gamma\,,\qquad&pp\to\bar{t}~{}(\to\ell^{-}\nu_{\ell}\bar{b})~{}\gamma\,,\\ pp\to t~{}(\to\ell^{+}\bar{\nu}_{\ell}b)~{}\bar{t}~{}(\to\bar{q}~{}\gamma)\,,\qquad&pp\to t~{}(\to q~{}\gamma)~{}\bar{t}~{}(\to\ell^{-}\nu_{\ell}\bar{b})\,,\end{split}$ (16)
@@
Signal and background processes involving top quarks are generated at LO using the five-flavour number scheme. For $V\gamma$+jets backgrounds, exclusive samples based on matrix elements including up to two additional partons are produced and matched to the parton shower using the MLM scheme Mangano:2006rw ; Alwall:2008qv , in order to avoid phase-space double counting. In addition, all processes employ a dynamical scale choice, defined as the average transverse mass of all massive final-state particles, and we use the NNPDF_40_lo_as_01180 set of parton distribution functions. In our simulation tool chain, the decays of all heavy resonances such as the top quark and electroweak gauge bosons are handled using MadSpin Artoisenet:2012st and MadWidth Alwall:2014bza to preserve spin correlations. The resulting parton-level events are then passed to Pythia 8.309 Bierlich:2022pfr for parton showering, hadronisation and hadron decays.
Detector effects are simulated using the Simplified Fast Simulation (SFS) module Araz:2020lnp ; Araz:2021akd embedded within the MadAnalysis 5 framework Conte:2012fm ; Conte:2014zja ; Conte:2018vmg , which is also used to implement the analysis strategy and prepare inputs for the machine learning models as introduced in Cornell:2024dki ; Cornell:2021gut . In this scheme, jets are reconstructed using the anti-$k_{T}$ algorithm Cacciari:2008gp with a radius parameter $R=0.4$ as implemented in FastJet 3.4.1 Cacciari:2011ma . Lepton and jet reconstruction and identification performance is taken from a validated MadAnalysis 5 routine designed for searches involving right-handed $W$ bosons and heavy neutrinos DVN/UMGIDL_2023 ; CMS:2021dzb , that has been described in detail in Frank:2023epx .
Finally, for the implementation of the machine learning models on which our analysis relies, we use the PyTorch Geometric platform fey2019fast to build and train graph neural networks, while the standard PyTorch package paszke2019pytorch is employed for fully connected networks and transformers. The Scikit-Learn library is used to facilitate preprocessing steps and evaluation metrics, whereas we rely on the PyTorch Lightning software falcon2019pytorch to manage the training workflows including model validation. For the purpose of open science, all code and analysis pipelines are available from a public GitHub repository.
@@ 4.1 Cut-based strategy
We initiate our cut-based analysis by applying a series of baseline preselection criteria to match the signal topology. Events are required to contain exactly one isolated charged lepton (either an electron or a muon) with transverse momentum $p_{T}>27$ GeV and pseudo-rapidity $|\eta|<2.5$, although for electrons we exclude the calorimeter transition regions of a typical LHC detector corresponding to $1.521.37<|\eta_{e}|<$ 1.52. Lepton isolation is enforced through the variable $I_{\ell}$, that we define as the scalar sum of the transverse momenta of all tracks $i$ separated by a transverse distance $\Delta R(i,\ell)<0.3$ from the lepton,
@@
$I_{\gamma}<\begin{cases}0.05\times p_{T}^{\gamma}\,\qquad\qquad\quad{\rm for}\qquad\Delta R(i,\gamma)<0.2\,,\\ 0.22\times p_{T}^{\gamma}+2.45\,\qquad{\rm for}\qquad\Delta R(i,\gamma)<0.4\,,\end{cases}$ (18)
@@
Jets are required to have $p_{T}>25$ GeV and $|\eta|<2.5$, this definition thereby rejecting forward jets, and events must contain at least one of such jet candidates. Among these jet candidates, we impose that at least one of them is $b$-tagged, which allows us to specifically target jets originating from the leptonic decay of top quarks. We remind that the detector response is parametrised following Frank:2023epx , thus using a $b$-tagging efficiency ($\mathcal{E}_{b\to b}$) and mistagging rates of a charm jet as a $b$-jet ($\mathcal{E}_{c\to b}$) and of a light jet as a $b$-jet ($\mathcal{E}_{j\to b}$) as a $b$-jet of
[…]
@@ 4.2.1 Generalities
In this section, we improve our analysis of top-photon FCNC interactions using three deep learning techniques: multi-layer perceptrons (MLPs), graph attention networks (GATs) and transformer-based networks. These networks have been chosen for their complementary strengths in handling the input data, which is structured as a set of features derived from the reconstructed kinematic properties of the final-state objects in each event. While MLPs operate on fixed-size vectors and treat all features independently, GATs and transformers can process richer and more complex event representations that preserve the relationships between the objects within an event.
@@ 4.2.2 Methodology
The learning process is iterative. In each training epoch defined as one complete pass over the dataset, the model updates its internal parameters to minimise a predefined loss function. Input events, represented as vectors, graphs or sets depending on the architecture, are fed into the network in mini-batches during the forward pass. Each input propagates through the network layers, and the output layer produces class scores that are converted into probabilities using a softmax activation function. To quantify the agreement between predicted probabilities and the true labels, we use a binary cross-entropy loss function where lower loss values indicate better predictive accuracy on the training data. After computing the loss, gradients are calculated via backpropagation, which applies the chain rule of calculus to evaluate how each weight and bias contributes to the final error. These gradients guide the update of the model parameters using stochastic gradient descent or an adaptive optimiser such as AdamW loshchilov2017decoupled . This optimisation cycle comprising forward propagation, loss evaluation, backpropagation and parameter update is repeated over multiple epochs until convergence. The trained model is then evaluated on an independent test set to assess its generalisation performance on unseen data.
@@
Feature embedding layers -- [512,256,128] FC [512,256,128]FC
@@
Number of attention heads -- 32 16
Drop out rate $5\%$ $25\%$ $25\%$
Pooling -- Average Average
Final MLP -- [128,64] FC [128,64] FC
@@
Finally, we remind that another source of variability in performance arises from fluctuations induced by the random partitioning of data into training, validation and test sets. To reduce this statistical variance and obtain a robust estimate of the model’s generalisation ability, we employ $k$-fold cross-validation. In this method, the dataset is divided into $k$ equal subsets. The model is then trained on $k{-}1$ of them and validated on the remaining one. This procedure is repeated $k$ times, with each subset serving once as the validation set. The final performance metric is then obtained by averaging over all $k$ runs. For all models in our study, we use $k=3$, ensuring consistent and statistically stable evaluation across architectures.
@@ 4.2.3 Multi-Layer Perceptron
Starting from a combination of low- and high-level kinematical observables, we employ an MLP to develop a classification strategy that outperforms the cut-based approach of section 4.1 by capturing characteristic patterns in both signal and background event samples. The fully connected structure of the MLP makes it then well suited for learning global correlations and translating them into a classification decision. In particular, the MLP is capable of identifying non-linear patterns that increase the separation power between signal and background Rumelhart:1986 ; Goodfellow:2016 .
@@
The adopted MLP architecture consists of an input layer with 29 neurons corresponding to the selected features. This is followed by four fully connected layers with 512, 256, 128 and 64 neurons respectively, each using a GeLU activation function. Dropout layers with a rate of 10% are placed after each hidden layer to mitigate overfitting, and the final layer is a single-neuron output with a sigmoid activation function providing a probability estimate for signal classification. For training, we use the AdamW optimiser loshchilov2017decoupled with a learning rate of $10^{-3}$ and a weight decay of $10^{-4}$ to minimise a binary cross entropy loss function.
@@
The classification performance of the MLP is summarised in figure 6 in which we show associated confusion matrices and receiver operating characteristic (ROC) curves. They both consist in valuable tools for evaluating classification models, though they serve different purposes and provide complementary insights. The confusion matrix offers a detailed, threshold-specific breakdown of predictions by reporting the number of true positives, true negatives, false positives and false negatives. This facilitates a granular understanding of the model’s performance in real-world decision-making scenarios, particularly in assessing the types of errors made. Derived metrics such as precision, recall and the F1-score are especially important when the misclassification costs are asymmetric Cornell:2021gut . By contrast, the ROC curve provides a threshold-independent evaluation by plotting the true positive rate against the false positive rate across all possible classification thresholds. It gives a global view of the model’s discriminative ability. However, a high ROC score given through the area under the curve (AUC) does not guarantee optimal performance at any specific threshold, where the model might still yield a high rate of false positives or miss true signals. Therefore, while the ROC curve summarises the overall classification potential, the confusion matrix remains essential for fine-tuning the model’s behaviour at specific operating points relevant to practical applications. In the top-left panel of figure 6, we show results for the three separate signal processes: $pp\to t\gamma$ and $pp\to\bar{t}t$ production followed by a rare $t\to u\gamma$ or $t\to c\gamma$ decay. The ROC curves indicate that the MLP achieves the highest discrimination for the single-top channel, with an AUC of 0.928. The $t\bar{t}$ production modes yield slightly lower but comparable performance, with AUCs of 0.789 and 0.778 for the charm and up-quark FCNC top decays respectively. The confusion matrices shown in the other three panels of the figure support these observations and confirm the consistent behaviour of the classifier across the different signal channels.
@@ 4.2.4 Graph Attention Network
Graph Neural Networks (GNNs) are powerful deep learning architectures designed to process graph-structured data. In the context of high-energy physics, such graphs often represent the relationships between final-state particles, with nodes corresponding to individual objects and edges encoding pairwise interactions or geometrical proximity. In classification tasks, GNNs aim to learn representations of signal and background events by aggregating information from each node’s neighbourhood. This aggregation captures both local graph structure and node features, which allows the model to distinguish between different classes of events. Among the various GNN architectures, Graph Attention Networks (GATs) have demonstrated particularly strong performance in particle physics applications Velickovic:2017lzs ; Yu:2023juh ; Calafiura:2024qhv ; Esmail:2023axd . Traditional GNNs use fixed or uniform weighting schemes during neighbourhood aggregation. However, this uniformity can limit the model’s ability to differentiate between relevant and irrelevant connections Shlomi:2020gdn ; Thais:2022iok ; Esmail:2024gdc ; Esmail:2024jdg ; Hammad:2025ewr . In contrast, GATs introduce an attention mechanism that dynamically assigns weights to neighbouring nodes during message passing. This enables the network to focus on the most informative parts of the graph for each node, which is particularly advantageous in heterogeneous or noisy environments where not all neighbours contribute equally. As a result, GATs offer improved interpretability and adaptability, often leading to higher classification accuracy and robustness in practical applications.
In GATs, the attention mechanism computes a pairwise attention score for each connected node pair, quantifying how much a given neighbour contributes to the target node’s feature update Esmail:2023axd . Specifically, the attention coefficient between node $i$ and its neighbour $j$ is defined as
@@
Figure 8 summarises the classification performance of the GAT network across the three considered processes: $pp\to t\gamma$, $pp\to\bar{t}t$ with a $t\to u\gamma$ decay and a $t\to c\gamma$ decay. The ROC curves reveal that the single top production process ($pp\to t\gamma$) achieves, like for our MLP implementation, the highest discrimination power with an AUC of $0.973$. Furthermore, the top pair production processes yield slightly lower AUC scores of $0.907$ and $0.926$ respectively, which is this time much larger than in the MLP case. Finally, we note that the confusion matrices reflect a consistent behaviour, showing a reliable separation of signal and background across all processes.

### v2

@@ 3.1 Modelling signal processes at the LHC
Figure 2: Leading-order cross sections for single top production in association with a photon via cgcg (or equivalently $\bar{c}g$) fusion (left), $\bar{u}g$ fusion (middle) and ugug fusion (right). The bands indicate theoretical uncertainties from scale and PDF variations summed in quadrature.
@@
as illustrated in figure 2 for a hadronic centre-of-mass energy $\sqrt{s}=13.6$ TeV. These predictions are obtained using Madgraph_aMC@NLO (v3.4.1) Alwall et al. (2014), with the NNPDF_40_lo_as_01180 PDF set Ball and others (2022) and renormalisation and factorisation scales set to the average transverse mass of the final-state particles, and an implementation of the relevant FCNC operators in FeynRules Christensen et al. (2011); Alloul et al. (2014) that has been exported in the UFO format Degrande et al. (2012); Darmé and others (2023). The resulting cross sections are found to span several orders of magnitude, from 1–10 fb for $\mathcal{O}(1)$ couplings down to 0.01–0.1 ab for couplings around $10^{-4}$ with the quadratic FCNC coupling dependence given in Eq. (11).
@@
This form for the cross section illustrates a clean separation between the QCD-driven $t\bar{t}$ production mechanism, one FCNC-induced top decay and one SM top decay. We compute the inclusive $t\bar{t}$ production cross section using Top++ v2.0 Czakon and Mitov (2014); Czakon et al. (2013), including next-to-next-to-leading-order (NNLO) QCD corrections matched to next-to-next-to-leading-logarithmic (NNLL) soft-gluon resummation. We use this time the NNPDF_40_nnlo_as_01180 PDF set, and we fix both the renormalisation and factorisation scales to the top mass.
@@ 3.2 Technical setup
[…]
@@
Signal and background processes involving top quarks are generated at LO using the five-flavour number scheme. For $V\gamma$+jets backgrounds, exclusive samples based on matrix elements including up to two additional partons are produced and matched to the parton shower using the MLM scheme Mangano et al. (2007); Alwall et al. (2009), in order to avoid phase-space double counting. In addition, all processes employ a dynamical scale choice, defined as the average transverse mass of all massive final-state particles, and we use the NNPDF_40_lo_as_01180 set of parton distribution functions. In our simulation tool chain, the decays of all heavy resonances such as the top quark and electroweak gauge bosons are handled using MadSpin Artoisenet et al. (2013) and MadWidth Alwall et al. (2015) to preserve spin correlations. The resulting parton-level events are then passed to Pythia 8.309 Bierlich and others (2022) for parton showering, hadronisation and hadron decays.
Detector effects are simulated using the Simplified Fast Simulation (SFS) module Araz et al. (2021, 2022) embedded within the MadAnalysis 5 framework Conte et al. (2013, 2014); Conte and Fuks (2018), which is also used to implement the analysis strategy and prepare inputs for the machine learning models as introduced in Cornell et al. (2025, 2022). In this scheme, jets are reconstructed using the anti-$k_{T}$ algorithm Cacciari et al. (2008) with a radius parameter $R=0.4$ as implemented in FastJet 3.4.1 Cacciari et al. (2012). Lepton and jet reconstruction and identification performance is taken from a validated MadAnalysis 5 routine designed for searches involving right-handed $W$ bosons and heavy neutrinos Jueid and Fuks (2023); Tumasyan and others (2022b), that has been described in detail in Frank et al. (2024b).
Finally, for the implementation of the machine learning models on which our analysis relies, we use the PyTorch Geometric platform Fey and Lenssen (2019) to build and train graph neural networks, while the standard PyTorch package Paszke et al. (2019) is employed for fully connected networks and transformers. The Scikit-Learn library is used to facilitate preprocessing steps and evaluation metrics, whereas we rely on the PyTorch Lightning software William Falcon and The PyTorch Lightning team (2019) to manage the training workflows including model validation. For the purpose of open science, all code and analysis pipelines are available from a public GitHub repository.
@@ 4.1 Cut-based strategy
We initiate our cut-based analysis by applying a series of baseline preselection criteria to match the signal topology. Events are required to contain exactly one isolated charged lepton (either an electron or a muon) with transverse momentum $p_{T}>27$ GeV and pseudo-rapidity $|\eta|<2.5$, although for electrons we exclude the calorimeter transition regions of a typical LHC detector corresponding to $1.37<|\eta_{e}|<1.52$. Lepton isolation is enforced through the variable $I_{\ell}$, that we define as the scalar sum of the transverse momenta of all tracks $i$ separated by a transverse distance $\Delta R(i,\ell)<0.3$ from the lepton,
@@
[…]
@@
Jets are required to have $p_{T}>25$ GeV and $|\eta|<2.5$, this definition thereby rejecting forward jets, and events must contain at least one of such jet candidates. Among these jet candidates, we impose that at least one of them is $b$-tagged, which allows us to specifically target jets originating from the leptonic decay of top quarks. We remind that the detector response is parametrised following Frank et al. (2024b), thus using a $b$-tagging efficiency ($\mathcal{E}_{b\to b}$) and mistagging rates of a charm jet as a $b$-jet ($\mathcal{E}_{c\to b}$) and of a light jet as a $b$-jet ($\mathcal{E}_{j\to b}$) as a $b$-jet of
$\begin{split}\mathcal{E}_{j\to b}=&\ 10^{-2}+3.8\times 10^{-5}\times p_{T}\,,\\ \mathcal{E}_{c\to b}=&\ 0.25\times\tanh(0.018\times p_{T})\times(1+0.0013\times p_{T})^{-1}\,,\\ \mathcal{E}_{b\to b}=&\ 0.85\times\tanh(0.0025\times p_{T})\times\frac{25}{1+0.063\times p_{T}}\,.\end{split}$ (19)
@@ 4.2.1 Generalities
In this section, we improve our analysis of top-photon FCNC interactions using three deep learning techniques: multi-layer perceptrons (MLPs), graph attention networks (GATs) and transformer-based networks [^1] In this work, signal and background samples simulated at leading order in QCD are used for both the training and validation stages. We stress out that higher-order predictions may induce shifts in the network latent space distribution due to impact on the kinematic distributions and their correlations. To mitigate such effects, the network architecture can be adapted using domain-adversarial techniques, which improve the generalisation of the latent space and enhance robustness against bias. This strategy has been discussed, for example, in Baalouch et al. (2019); Bishara et al. (2024). Whereas advanced architectures with a large number of tunable parameters, such as graph neural networks and transformers, are particularly well suited to incorporate this and address the issue effectively, such an explorative study is left for future work.. These networks have been chosen for their complementary strengths in handling the input data, which is structured as a set of features derived from the reconstructed kinematic properties of the final-state objects in each event. While MLPs operate on fixed-size vectors and treat all features independently, GATs and transformers can process richer and more complex event representations that preserve the relationships between the objects within an event.
@@ 4.2.2 Methodology
The learning process is iterative. In each training epoch defined as one complete pass over the dataset, the model updates its internal parameters to minimise a predefined loss function. Input events, represented as vectors, graphs or sets depending on the architecture, are fed into the network in mini-batches during the forward pass. Each input propagates through the network layers, and the output layer produces class scores that are converted into probabilities using a softmax activation function. To quantify the agreement between predicted probabilities and the true labels, we use a binary cross-entropy loss function where lower loss values indicate better predictive accuracy on the training data. After computing the loss, gradients are calculated via backpropagation, which applies the chain rule of calculus to evaluate how each weight and bias contributes to the final error. These gradients guide the update of the model parameters using stochastic gradient descent or an adaptive optimiser such as AdamW Loshchilov and Hutter (2017). This optimisation cycle comprising forward propagation, loss evaluation, backpropagation and parameter update is repeated over multiple epochs until convergence. The trained model is then evaluated on an independent test set to assess its generalisation performance on unseen data.
@@
Feature embedding layers - [512,256,128] FC [512,256,128]FC
@@
Number of attention heads - 32 16
Drop out rate $5\%$ $\% 2525\%$
Pooling - Average Average
Final MLP - [128,64] FC [128,64] FC
@@
Finally, we remind that another source of variability in performance arises from fluctuations induced by the random partitioning of data into training, validation and test sets. To reduce this statistical variance and obtain a robust estimate of the model’s generalisation ability, we employ $k$-fold cross-validation. In this method, the dataset is divided into $k$ equal subsets. The model is then trained on k1k{-}1 of them and validated on the remaining one. This procedure is repeated $k$ times, with each subset serving once as the validation set. The final performance metric is then obtained by averaging over all $k$ runs. For all models in our study, we use $k=3$, ensuring consistent and statistically stable evaluation across architectures.
@@ 4.2.3 Multi-Layer Perceptron
Starting from a combination of low- and high-level kinematical observables, we employ an MLP to develop a classification strategy that outperforms the cut-based approach of section 4.1 by capturing characteristic patterns in both signal and background event samples. The fully connected structure of the MLP makes it then well suited for learning global correlations and translating them into a classification decision. In particular, the MLP is capable of identifying non-linear patterns that increase the separation power between signal and background Rumelhart et al. (1986); Goodfellow et al. (2016).
@@
The adopted MLP architecture consists of an input layer with 29 neurons corresponding to the selected features. This is followed by four fully connected layers with 512, 256, 128 and 64 neurons respectively, each using a GeLU activation function. Dropout layers with a rate of 10% are placed after each hidden layer to mitigate overfitting, and the final layer is a single-neuron output with a sigmoid activation function providing a probability estimate for signal classification. For training, we use the AdamW optimiser Loshchilov and Hutter (2017) with a learning rate of $10^{-3}$ and a weight decay of $10^{-4}$ to minimise a binary cross entropy loss function.
@@
The classification performance of the MLP is summarised in figure 6 in which we show associated confusion matrices and receiver operating characteristic (ROC) curves. They both consist in valuable tools for evaluating classification models, though they serve different purposes and provide complementary insights. The confusion matrix offers a detailed, threshold-specific breakdown of predictions by reporting the number of true positives, true negatives, false positives and false negatives. This facilitates a granular understanding of the model’s performance in real-world decision-making scenarios, particularly in assessing the types of errors made. Derived metrics such as precision, recall and the F1-score are especially important when the misclassification costs are asymmetric Cornell et al. (2022). By contrast, the ROC curve provides a threshold-independent evaluation by plotting the true positive rate against the false positive rate across all possible classification thresholds. It gives a global view of the model’s discriminative ability. However, a high ROC score given through the area under the curve (AUC) does not guarantee optimal performance at any specific threshold, where the model might still yield a high rate of false positives or miss true signals. Therefore, while the ROC curve summarises the overall classification potential, the confusion matrix remains essential for fine-tuning the model’s behaviour at specific operating points relevant to practical applications. In the top-left panel of figure 6, we show results for the three separate signal processes: $pp\to t\gamma$ and $pp\to\bar{t}t$ production followed by a rare $t\to u\gamma$ or $t\to c\gamma$ decay. The ROC curves indicate that the MLP achieves the highest discrimination for the single-top channel, with an AUC of 0.928. The $t\bar{t}$ production modes yield slightly lower but comparable performance, with AUCs of 0.789 and 0.778 for the charm and up-quark FCNC top decays respectively. The confusion matrices shown in the other three panels of the figure support these observations and confirm the consistent behaviour of the classifier across the different signal channels.
@@ 4.2.4 Graph Attention Network
Graph Neural Networks (GNNs) are powerful deep learning architectures designed to process graph-structured data. In the context of high-energy physics, such graphs often represent the relationships between final-state particles, with nodes corresponding to individual objects and edges encoding pairwise interactions or geometrical proximity. In classification tasks, GNNs aim to learn representations of signal and background events by aggregating information from each node’s neighbourhood. This aggregation captures both local graph structure and node features, which allows the model to distinguish between different classes of events. Among the various GNN architectures, Graph Attention Networks (GATs) have demonstrated particularly strong performance in particle physics applications Veličković et al. (2017); Yu et al. (2023); Calafiura et al. (2024); Esmail et al. (2023). Traditional GNNs use fixed or uniform weighting schemes during neighbourhood aggregation. However, this uniformity can limit the model’s ability to differentiate between relevant and irrelevant connections Shlomi et al. (2020); Thais et al. (2022); Esmail et al. (2024b, a); Hammad and Jueid (2025). In contrast, GATs introduce an attention mechanism that dynamically assigns weights to neighbouring nodes during message passing. This enables the network to focus on the most informative parts of the graph for each node, which is particularly advantageous in heterogeneous or noisy environments where not all neighbours contribute equally. As a result, GATs offer improved interpretability and adaptability, often leading to higher classification accuracy and robustness in practical applications.
In GATs, the attention mechanism computes a pairwise attention score for each connected node pair, quantifying how much a given neighbour contributes to the target node’s feature update Esmail et al. (2023). Specifically, the attention coefficient between node $i$ and its neighbour $j$ is defined as
@@
Figure 8 summarises the classification performance of the GAT network across the three considered processes: $p\to t\gamma, pp$$p\to\bar{t}t$ with a $t\to u\gamma$ decay and a $t\to c\gamma$ decay. The ROC curves reveal that the single top production process ($pp\to t\gamma$) achieves, like for our MLP implementation, the highest discrimination power with an AUC of $0.973$. Furthermore, the top pair production processes yield slightly lower AUC scores of $0.907$ and $0.926$ respectively, which is this time much larger than in the MLP case. Finally, we note that the confusion matrices reflect a consistent behaviour, showing a reliable separation of signal and background across all processes.